Find the Wronskian of the following set of function
{sin 3x,cos 3x}
1
Expert's answer
2021-09-22T00:42:35-0400

We consider "y_1=\sin 3x" and "y_2=\cos 3x" . Then the requested wronskian is:


"W(y_1,y_2)=\begin{vmatrix}
y_1 & y_2\\
y_1' & y_2'
\end{vmatrix}=y_1y_2'-y_1'y_2"

"\left. 
\begin{array}{l}
y_1=\sin 3x\Rightarrow y_1'=3\cos 3x\\
y_2=\cos 3x\Rightarrow y_2'=-3\sin 3x
\end{array}
\right\}"

Substituting the latter into the Wronskian, we obtain:


"W(\sin 3x, \cos 3x)=\begin{vmatrix}
\sin 3x & \cos 3x\\
3\cos 3x & -3\sin 3x
\end{vmatrix}=-3\sin^23x+3\cos^2 3x="

"=3(\cos^2 3x-\sin^2 3x)=3\cos (2\cdot 3x)=\boxed{3\cos 6x}"